### v1

@@ 5.2 JAX framework
In this section we compare the performance of JAX with all other DP-SGD frameworks (all of them are based on PyTorch). To make a fair comparison between frameworks, we implemented Poisson subsampling and virtual batching, based on the Opacus implementation.
@@
Compilation time Comparing JAX to PyTorch requires taking the compilation time into account that the DP-SGD implementations in PyTorch do not utilize. There is no straightforward way of calculating the compilation time, but we measure it as the duration to process the first physical batch. The execution times for each batch shows that the first one takes much more time than the others, therefore including the compilation time (Figure A.1). To provide a fair comparison, we also implemented a non-private JAX training using the same virtual batching as PyTorch.
@@
Throughput comparision In Figure 6, both JAX implementations start at the same point, since by using virtual batching, they are processing just one sample at a time and getting the gradients from the compiled function. The non-private benefits more from a larger batch size, although its throughput decreases at 128 instances. The compilation time (see Figure A.1) grows with the batch size. For the private model, it takes more time since the compiled function is more complex than the non-private counterpart. It includes expanding the dimensions and clipping the gradients, while the non-private directly computes the gradient of the whole mini-batch.
@@
JAX per-example clipping reaches a higher physical batch size, and the throughput is always higher than its Opacus counterpart. However, the Book Keeping Ghost Clipping implementation is closer to the JAX private version. They are the same until the physical batch size of 32, but even then, the variance in the JAX execution is high, meaning that it can even be below the BK-ghost throughput. JAX private implementation throughput drops when the batch size doubles to 64. This comparison would be a reason to use Book Keeping ghost clipping for Vision Transformer models over JAX implementation. Its throughput is comparable to the execution of compiled code from JAX while reaching a larger batch size.
JAX private model starts with a similar throughput as the non-private model and then steadily grows until it surpasses the Torch throughput after a physical batch size of 16. After that, JAX will always have better throughput, except at a physical batch size of 128, where JAX throughput drops and is the same as the PyTorch non-private. We ran the experiment multiple times, and at the 128 size, it has a low variance, with always the same throughput value. Another difference between the two frameworks is the variability in the experiments. PyTorch runs are remarkably consistent in having a low variance, and the same throughput result is expected every time for a fixed seed.
@@
JAX executions are more variable than those of PyTorch, likely due to its sensitivity to HPC environment fluctuations and accelerator stochasticity, as noted in Figure A.1. Another contributing factor is JAX’s asynchronous dispatch method, which complicates time benchmarking by issuing a promise rather than immediate results, concealing Python overheads.
@@
Poisson sampling Using JAX for DP introduces complexities, particularly around subsampling which is crucial for privacy accounting. Implementing Poisson subsampling results in variable batch sizes in JAX; changes in batch size require JIT to recompile, leading to graph retracing which is costly and contributes to execution run variability, as discussed by Chua et al. (2024).
@@
Comparison with PyTorch Although compiling PyTorch is possible, we could not see any improvements in terms of speed-up. While compiling the non-private model worked, the speed-up gained was minimal and, in the end, even lower if we consider the compilation time. PyTorch also recompiles after a batch size change. While trying to compile, PyTorch falls back to predefined CUDA operations that are already optimized. In the case of the private setting, the compilation does not recognize Opacus hooks and continues the execution without compiling them.
@@
JAX reaches a higher accuracy in Table A4 than its PyTorch counterpart. We are still evaluating what is causing the significant change, even when using the same hyperparameters and training for just two epochs. The difference is more prominent for the non-private case, but we still see the DP utility cost.
5.3 Lower precision
We consider using lower precision to speed up computational times. We evaluate the use of Tensor Core 32 (TF32) for training. It has 10 bits for precision, with eight range bits, the same as 32 single precision (FP32)Kharya (2020). Using lower precision can have benefits exactly where DP training struggles: it requires less memory, uses fewer bits to represent the data, and its operations are optimized for GPU, making them much faster NVIDIA (2023). It is specially optimized for the A100 GPU and unavailable for the V100, so the comparison will be between training on the A100 with and without TF32.
Experimental results In Figure 7, the relative difference between mean throughput between runs with TF32 and FP32. For non-private training, throughput increases with model size, peaking at nearly four times the speed for the ViT Huge model. Then, the private training throughput increases for the smaller models, but it goes down again as the model size grows after the base size. The ViT Base model presents the optimal speed-up, suggesting that the best throughput can be obtained using a middle-size model. One that is too small does not gain much from TF32, and the larger ones are too expensive. If we compare the non-private, the growth between the large and huge model sizes is minimal. The model size is the bottleneck, and the private one is the most affected by it due to the nature of the clipping algorithm. However, this remains open for further investigation on why the private training goes up and down with the model size in lower precision. Regarding the memory advantages by TF32, we could not see an improvement. Both models, with and without TF32, could fit the same number of instances.
Concerns regarding TF-32 under DP There are two concerns with using lower precision in DP deep learning: its effects on utility and privacy. For the first issue, using lower precision may affect utility, as it is less precise. We did not find a significant decay in the accuracy of the models; it differs by decimal points at the precision (See Table A4 in the Appendix). For example, private training using the optimal number of training steps, both models reach the same accuracy of with a significant speed-up thanks to lower precision.
The second concern about privacy was first highlighted by Mironov (2012), but no fully satisfactory solution exists yet. The problem is that theoretical guarantees for most algorithms assume computations are performed on reals, and finite precision arithmetic can lead to actual violations. Discrete mechanisms (e.g. Canonne et al., 2020; Agarwal et al., 2021) that avoid the theoretical challenges exist, but they are often less convenient and may lead to loss of utility, especially in low precision settings. The efficiency of different discrete mechanisms in TF-32 is an interesting topic of further research.
@@ 6 Multi-GPU Training
6.1 Experimental Setup
We utilize V100 GPUs on HPC nodes that have 4 GPUs per node. The other experimental setting is idential to the one in Section 4. Results for utilizing up to 24 A100 GPUs can be found in the Appendix Figure A.3.
We focus on comparing the scaling behaivour between the non-private baseline that uses PyTorch and the DP-SGD implementation using Opacus. Both frameworks provide mature tooling for distributed training. Although the other clipping methods do not break, we are still determining if they are correctly handling the distributed gradient computation and clipping or if distributed training is even supported.
6.2 Experimental Results
Figure 8 shows the throughput increase as we add more computational resources, going from one to 80 GPUs. The plot shows that the throughput does not grow linearly and starts changing from the ideal after using more than one node (i.e. when using more than 4 GPUs). The communication inside the node is fast, but the communication between nodes will always be slower. The bottleneck is the bandwidth, and it prevents the model from scaling linearly. Notably, it affects the non-private training baseline much more, while the private scales better. For the 80 GPUs, the private training achieves of the ideal throughput, and the non-private training only achieves . Private training scales better because it is slower and only sometimes saturates the network with updates.
@@ 7 Conclusion
While DP-SGD is significantly more costly than non-private training, we identified feasible speed-ups that are often easy to apply but have some drawbacks. These are: 1 More efficient implementations of DP-SGD which additionally decrease the memory footprint (allowing for training larger models). However, the implementations are not as mature as Opacus and do not support all types of neural network layers (yet). 2 JAX which processes samples faster than PyTorch but looses the advantage through frequent re-compilations when utilizing proper Poisson sampling, does not offer a comprehensive DP-SGD implementation as PyTorch and runs not as stable as PyTorch. 3 Lower Precision using TF-32 which increases the throughput but the implications on the theoretical guarantees of DP-SGD need to be explored in future work. Finally, we found that distributed computing using DP-SGD scales better than non-private training and allows for fast training of models.
Acknowledgements
This work was supported by the Finnish Ministry of Education and Culture and CSC - IT Centre for Science (Decision diary number OKM/10/524/2022), the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Niki Loppi contributed under the NVIDIA AI Technology Center (NVAITC) Finland program. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work has been performed using resources provided by the CSC – IT Center for Science, Finland. We thank Joonas Jälkö for helpful discussions regarding implementing DP-SGD with JAX.
@@ A.2 Models
Vision Transformer (ViT) Dosovitskiy et al. (2021). Taken from https://huggingface.co/timm/vit_base_patch16_224.augreg2_in21k_ft_in1k
Big Transfer BiT-ResNet Kolesnikov et al. (2020). Taken from https://github.com/google-research/big_transfer
@@ A.3 Hyperparameters
We use the hyperparameters obtained on request from Tobaben et al. (2023). The hyperparameters for both models are in Table A2. We do not optimize them futher, as model utility is not the main objective in this work.
@@
| BiT-ResNet | All | 8 | 0.00098 | 6.53 |
@@ Appendix B Additional Results
| Section | PyTorch non-private | Opacus per-example |
| Forward | 81.14 | 101.53 |
| Backward | 163.85 | 681.48 |
| Clip and Accumulation | 0 | 26.76 |
| Optimizer Step | 38.17 | 99.65 |
@@
| O-flat Opacus | 0.7096 |
| O-flat Opacus/TF32 | 0.7094 |
| O-flat JAX | 0.8009 |
| PV-ghost | 0.6978 |
| BK-ghost | 0.7341 |

### v2

@@ 6 JAX
In this section we compare the performance of the two JAX implementations with all other DP-SGD frameworks (all of them are based on PyTorch).
@@
Compilation time Comparing JAX to PyTorch requires taking the compilation time into account that the DP-SGD implementations in PyTorch do not utilize. There is no straightforward way of calculating the compilation time, but we measure it as the duration to process the first batch. The execution times for each batch shows that the first one takes much more time than the others, which means it includes the compilation time (Figure A.2). To provide a fair comparison, we also implemented a non-private JAX training using the same virtual batching as PyTorch.
@@
Throughput comparison In Figure 6, both naive JAX implementations (private and non-private) start at the same point, since by using virtual batching, they are processing just one sample at a time and getting the gradients from the compiled function. The non-private benefits more from a larger batch size, although its throughput decreases at 128 instances. The compilation time (see Figure A.2) grows with the batch size. For the private model, it takes more time since the compiled function is more complex than the non-private counterpart. It includes expanding the dimensions and clipping the gradients, while the non-private directly computes the gradient of the whole mini-batch.
The masked DP-SGD jax implementation shows a higher throughput than any other private implementations. Since for this implementation we are fitting the whole logical batch in CPU memory, we can split it and have static sizes. Therefore the compilation time will be higher for the first logical batch, but for the next iterations we can see the gains in speed, as it does not need to recompile. Its throughput changes less with respect to its batch size, in comparison to other methods. It resembles an efficient DP-SGD implementation in JAX using proper Poisson subsampling.
@@
JAX per-example clipping reaches a higher physical batch size, and the throughput is always higher than its Opacus counterpart. However, the Book Keeping Ghost Clipping implementation is closer to the JAX private version. They are the same until the physical batch size of 32, but even then, the variance in the JAX execution is high, meaning that it can even be below the BK-ghost throughput. JAX private naive implementation throughput drops when the batch size doubles to 64. For practitioners that remain in PyTorch, Book Keeping ghost clipping presents a throughput comparable to the execution of the naive JAX while reaching a larger batch size.
@@
Another difference between the two frameworks is the variability in the experiments. PyTorch runs are remarkably consistent in having a low variance, and the same throughput result is expected every time for a fixed seed. JAX executions are more variable than those of PyTorch, likely due to its sensitivity to HPC environment fluctuations and accelerator stochasticity, as noted in Figure A.2. Another contributing factor is JAX’s asynchronous dispatch method, which complicates time benchmarking by issuing a promise rather than immediate results, concealing Python overheads Lastly, the naive method recompiles everytime it process a batch with a different size, adding to the time variability.
@@
Poisson sampling Using JAX for DP introduces complexities, particularly around subsampling which is crucial for privacy accounting. Implementing Poisson subsampling results in variable batch sizes in JAX; changes in batch size require JIT to recompile, leading to graph retracing which is costly and contributes to execution run variability, as discussed by Chua et al. (2024). Our masked DP-SGD implementation overcomes this issue while using proper Poisson subsampling and therefore ensuring the correct privacy budget.
@@
PyTorch Compilation Although compiling PyTorch is possible, we could not see any improvements in terms of speed-up. While compiling the non-private model worked, the speed-up gained was minimal and, in the end, even lower if we consider the compilation time. PyTorch also recompiles after a batch size change. While trying to compile, PyTorch falls back to predefined CUDA operations that are already optimized. In the case of the private setting, the compilation does not recognize Opacus hooks and continues the execution without compiling them.
@@ 7 Multi-GPU Training
We utilize V100 GPUs on HPC nodes that have 4 GPUs per node. The other experimental setting is identical to the one in Section 4. Results for utilizing up to 24 A100 GPUs can be found in Figure A.4 in the Appendix. We focus on comparing the scaling behaivour between the non-private baseline that uses PyTorch and the DP-SGD implementation using Opacus. Both frameworks provide mature tooling for distributed training.
Figure 7 shows the throughput increase as a function of number of GPUs. The throughput does not grow linearly and starts changing from the ideal linear scaling after using more than one node (i.e. when using more than 4 GPUs). The communication inside the node is fast, but the communication between nodes will always be slower. The bottleneck is the bandwidth, and it prevents the model from scaling linearly. Notably, it affects the non-private training baseline much more, while the private scales close to optimal up to 32 GPUs. For the 80 GPUs, the private training achieves of the ideal throughput, and the non-private training only achieves . Private training scales better because it is slower and only sometimes saturates the network with updates.
@@ 8 Conclusion
While DP-SGD is significantly more costly than non-private training, we identified feasible speed-ups that are often easy to apply but have some drawbacks. These are: 1 More efficient implementations of DP-SGD which additionally decrease the memory footprint (allowing for training larger models). However, the implementations are not as mature as Opacus and do not support all types of neural network layers (yet). 2 JAX which processes samples faster than PyTorch but looses the advantage through frequent re-compilations when utilizing proper Poisson sampling, does not offer a comprehensive DP-SGD implementation as PyTorch and exhibits a greater variability in execution times than PyTorch. 3 We present an efficient implementation DP-SGD with JAX that complies with the Poisson sampling requirements while using physical batches of the same length, also complying with JAX efficient optimizations. 4 Lower Precision using TF32 which increases the throughput but the implications on the theoretical guarantees of DP-SGD need to be explored in future work. Finally, we found that distributed computing using DP-SGD scales better than non-private training and allows for fast training of models.
Acknowledgments
This work was supported by the Finnish Ministry of Education and Culture and CSC - IT Centre for Science (Decision diary number OKM/10/524/2022), the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Niki Loppi contributed under the NVIDIA AI Technology Center (NVAITC) Finland program. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work has been performed using resources provided by the CSC – IT Center for Science, Finland and the Finnish Computing Competence Infrastructure (FCCI). We thank Talal Alrawajfeh for helpful discussions regarding implementing DP-SGD with JAX.
@@ A.2 Models
Vision Transformer (ViT) (Dosovitskiy et al., 2021). Taken from https://huggingface.co/timm/vit_base_patch16_224.augreg2_in21k_ft_in1k
Big Transfer ResNet (Kolesnikov et al., 2020). Taken from https://github.com/google-research/big_transfer
@@ A.3 Hyperparameters
We use the hyperparameters obtained on request from Tobaben et al. (2023). The hyperparameters for both models are in Table A2. Even though model utility is not the main objective in this work, in the non-private case, the learning rate is suboptimal. By changing it to 0.00027 we see an accuracy improvement, therefore the one we are using.
@@
| ResNet | All | 8 | 0.00098 | 6.53 |
A.4 Poisson Subsampling JAX Algorithm
We present our DP-SGD implementation in JAX that uses the correct Poisson subsampling and therefore we can account for its privacy. The main problem with implementing DP-SGD with JAX is the batches of variable size. In order to address this issue, we compute always full physical batches and mask out gradients so that the total number of used gradients is equal the sampled logical batch sizes. This means that we always compute a little more gradients that required due to sampling. This prevents the recompiling.
@@ Appendix B Additional Results
| Opacus | 0.8732 |
| Opacus/TF32 | 0.8746 |
| JAX | 0.8009 |
| PV-ghost | 0.8752 |
| BK-ghost | 0.8752 |